# Continuous Test-time Domain Adaptation for Efficient Fault Detection under Evolving Operating Conditions

### v1

@@ Abstract
Fault detection is essential in complex industrial systems to prevent failures and optimize performance by distinguishing abnormal from normal operating conditions. With the growing availability of condition monitoring data, data-driven approaches have seen increased adoption in detecting system faults. However, these methods typically require large, diverse, and representative training datasets that capture the full range of operating scenarios, an assumption rarely met in practice, particularly in the early stages of deployment. Industrial systems often operate under highly variable and evolving conditions, making it difficult to collect comprehensive training data. This variability results in a distribution shift between training and testing data, as future operating conditions may diverge from previously observed ones. Such domain shifts hinder the generalization of traditional models, limiting their ability to transfer knowledge across time and system instances, ultimately leading to performance degradation in practical deployments. To address these challenges, we propose a novel method for continuous test-time domain adaptation, designed to support robust early-stage fault detection in the presence of domain shifts and limited representativeness of training data. Our proposed framework –Test-time domain Adaptation for Robust fault Detection (TARD) – explicitly separates input features into system parameters and sensor measurements. It employs a dedicated domain adaptation module to adapt to each input type using different strategies, enabling more targeted and effective adaptation to evolving operating conditions. We validate our approach on two real-world case studies from multi-phase flow facilities, delivering substantial improvements over existing domain adaptation methods in both fault detection accuracy and model robustness under real-world variability.
@@
This work proposes a novel method for efficient fault detection that continuously adapts to evolving operating conditions during test-time.
@@ 1 Introduction
To address this data scarcity, one strategy is to leverage information from ”experienced” fleet units – systems with extensive operational histories and rich datasets. A viable solution is to directly apply models trained on these well-instrumented units to less-monitored or newly deployed ones. This cross-unit deployment aims to enhance fault detection performance across the fleet, particularly for units with limited training data. However, such model transfer is hindered by a fundamental limitation: data-driven models commonly rely on the assumption that training and testing data are independently and identically distributed (i.i.d) [6]. In reality, industrial systems operate under diverse and dynamically changing conditions, violating this assumption. As a result, significant distribution shifts often exist between different units in the fleet. A model trained on one unit may therefore fail to generalize to another, frequently resulting in high false alarm rates and reduced detection accuracy. These limitations ultimately undermine the ability to fully capitalize on shared fleet knowledge.
@@
Domain adaptation (DA) has been extensively explored to address distribution shifts in data-driven models [7, 8]. These methods aim to align a labeled source domain with a related, unlabeled target domain. However, fault detection in industrial systems presents unique challenges. First, DA approaches typically assume the availability of labeled data in the source domain, which is especially scarce or unavailable under faulty conditions in both source domain and target domain. As a result, many commonly used unsupervised DA methods – which rely on source domain labels – are not applicable to fault detection tasks in industrial settings. Second, most DA methods treat domain shift as a single, static distributional difference between discrete domains. However, in complex industrial systems, the operating conditions evolve continuously, leading to domain shifts even within the same unit over time. Therefore, domain adaptation should go beyond discrete domain transfers and be capable of handling continuously evolving shifts. Relying on a fixed, discrete target domain is insufficient for ensuring robust fault detection in such dynamic environments.
@@
To address these challenges, we propose Test-time domain Adaptation for Robust fault Detection (TARD), a novel framework for unsupervised, continual test-time domain adaptation for fault detection that operates entirely without labeled data in the source domain. TARD enables robust fault detection across different units in industrial systems by continuously adapting to dynamically changing operating conditions during test time.
@@ 2.1 Fault Detection
Given this limitation, the common approach is to first shift the objective to detecting any form of abnormal behavior in the system. This process begins with unsupervised anomaly detection, which involves learning a model of normal system operation and the identifying any data patterns that deviates from the baseline [10]. These detected anomalies indicates potential problems, which can be analyzed further to determine if they represent a fault that could lead to a failure of the system [11, 9, 12].
Those unsupervised learning methods can be broadly categorized into three main categories: probabilistic models, one-class classification methods, and reconstruction-based approaches [13]. Probabilistic models aim to approximate the probability distribution of normal data, using the estimated density or an associated score function to assign anomaly scores, where lower values typically indicate more anomalous behavior. Various probabilistic models have been applied for this purpose, ranging from classic density estimation models (e.g. gaussian mixture models) [14] to energy-based models (EBMs) [15], predicting anomalies through estimation of the normal data probability distribution. One-class classification models take a different approach by learning a discriminative decision boundary that encloses the region of normal data, without explicitly estimating the full data distribution [13, 16, 17, 18]. These models aim to learn a compact representation of normal behavior and flag data sample that significantly deviates from the normal data distribution as an anomaly[19, 20]. Reconstruction-based methods, such as autoencoders (AEs), are trained to accurately reconstruct normal data samples. Anomalies are then identified based on reconstruction errors –when a test input deviates significantly from the learned normal patterns, the model struggles to reconstruct it, resulting in a high error that signals a potential fault [21, 22]. These models are expected to fit the data distribution under healthy conditions and then raise an alarm for predictions with large deviations when the test data distribution is significantly different from the learned distribution. Despite their usefulness, these unsupervised approaches require large, diverse, and representative datasets that comprehensively cover normal operating conditions, which is rarely fulfilled during the early operational phase of newly deployed systems, making robust early fault detection particularly challenging.
@@ 2.3 Domain Adaptation Applied to Fault Detection
However, most existing DA techniques rely on strong assumptions. They typically presume that abundant source and target samples are available to capture the underlying data distributions, an assumption that rarely holds in industrial settings, particularly for newly installed systems where only limited data can be collected, impeding timely condition monitoring [35]. Access to source data may also be restricted due to privacy or proprietary concerns, making unsupervised domain adaptation unfeasible [1].
Methods such as the one proposed in [35] adopt a one-class classification approach to address the challenge of limited data samples. However, they are generally designed for adaptation to one or more discrete, static target domains. In reality, industrial systems often operate under continuously evolving conditions due to environmental fluctuations, component aging, and ongoing system upgrades, leading to non-stationary distributions that static DA models cannot adequately address [37]. As a result, domain adaptation strategies tailored for discrete or static domains are ill-suited to dynamic industrial environments, where continuous, on-the-fly adaptation to domain shifts is required.
Test-time adaptation (TTA) has emerged as a promising approach to address the aforementioned limitations. TTA methods adapt pre-trained models during inference by leveraging only the distribution of the current test batch distribution [38, 39]. By dynamically updating model statistics or minimizing prediction entropy at inference, TTA enables real-time adaptation to distribution shifts, making it particularly attractive for PHM applications in environments with continuously evolving operating conditions [40].
@@
Existing methods tend to address isolated aspects of these challenges, but rarely in combination. Fleet-based approaches enable newly deployed units to leverage fleet-wide knowledge, but often fail to adapt to the significant variability of operating conditions within the fleet itself. While DA methods have gained traction, most cannot handle continuous domain shifts and depend on access to sufficiently representative source and target data. Moreover, TTA in the context of unsupervised detection risks model overfitting due to the absence of supervised faulty data. To the best of our knowledge, there is currently no approach that enables unsupervised fault detection under dynamically changing, data-scarce environments.
@@ 3.1 Problem Definition
The primary motivation for this research is to enable the transfer of operating experience across systems within a fleet – specifically, from systems with abundant monitoring data to other systems operating with the adaptation ability to evolving operating conditions. A key challenge addressed in this work is complete absence of labeled fault data in both the source and target domains. We assume that only healthy operating data is available and that neither the source nor target systems have experienced any faults during the training period. Additionally, we focus on target systems that are newly taken into operation, where only a limited number of observations can be collected for training. Moreover, the data distribution for these systems can evolve continuously due to changes in operating conditions and environmental factors. The objective is to adapt a prediction model trained on the original system so that it can provide accurate predictions for new systems and fleets, even when only a small amount of training data is available. Given:
@@ 3.2 System Variables
In industrial systems, input data typically comprise a mixture of variables reflecting different aspects of system behavior. To effectively handle domain shifts while preserving anomaly detection performance, it is important to distinguish between variables that represents changes of operating conditions or responses of system operating states. Therefore, we define two groups within the input parameters: $X=[x,w]$. $w$ denotes control variables, indicating variables that control system conditions. These variables are set by the operators or by the control system to optimize the performance under specified conditions. $x$ represents sensor measurements, which are sensor signals monitoring system components and reflecting real-time system states. Here, we assume that changes in the distribution of control variables do not necessarily indicate an abnormal status but rather distinct operating conditions to which we should adapt.
@@ 3.3 Reconstruction-based Anomaly Detection Framework
We deploy the paradigm of reconstruction-based pipeline for unsupervised anomaly detectopn as depicted in Figure 1. This method assumes that a model trained exclusively on normal (healthy) data can learn to reconstruct typical system behavior and deviations from this behavior with large reconstruction errors can be interpreted as potential anomalies. We deploy an autoencoder (AE), denoted as $f_{\theta}$, as our reconstruction model. It is trained exclusively on normal source data samples $X_{s}$ to accurately model the normal data distribution with accurately predicted signal value $\hat{X_{s}}$. The training objective is to minimize the mean-squared error (MSE) between the original data samples $X_{s}$, and their reconstructed counterparts $\hat{X_{s}}$:
@@ 4 Methodology
Figure 2: Test-time domain Adaptation for Robust fault Detection (TARD). a) We first pretrain the reconstruction-based anomaly detection model on the source dataset; b) For domain adaptative prediction, we add domain adaptive module and train it on the target training data.
@@ 4.2 Test-time Domain Adaptation Anomaly Detection
The core of the TARD framework is its ability to adapt to changing operating conditions during inference. As discussed in Section 3.2, control variables $w$ represent operational inputs that are manually or automatically adjusted during system operation. These variables indicate intentional changes in the system’s configuration and are not themselves indicators of abnormal behavior. Sensor measurements $x$, in contrast, reflect the real-time state of the system and are the primary source for fault detection. Therefore, as shown in Figure 2(a), for the reconstruction-based anomaly detection model, control variables are included alongside sensor measurements as part of the input to the autoencoder while the output only consists of the sensor measurements $x$. By excluding control variables from the reconstruction target and subsequent anomaly scoring, the model reduces the likelihood of false alarms triggered by expected changes in operating conditions.
@@
Based on the modified reconstruction framework introduced, we integrate an addition adaptive module, $h_{\phi}$, for test-time domain adaptation to bridge the domain gap between the source and target domains. The decision to incorporate a separate adaptive module, $h_{\phi}$, rather than embedding adaptive layers directly into the reconstruction model, stems from limitations observed in unsupervised anomaly detection. TTA methods, such as AdaBN, adapt to each batch during test time, inevitably fitting the distribution of abnormal data points. This impairs the model’s ability to distinguish between normal and abnormal samples. Instead, our adaptive module takes the predicted value and original controlled system variables as inputs. By excluding monitoring data signals from adaptation, this approach prevents overfitting to the potentially faulty data distribution, thereby preserving the model’s capability to accurately distinguish anomalies.
@@
During the adaptation phase, the pre-trained autoencoder $f_{\theta}$ is frozen, and the adaptive module $h_{\phi}$ is trained on a few target data samples to predict $\Delta x$, aimed at compensating for the large prediction errors due to the domain gap between source and target data. This adaptive module exclusively processes the control variables $w$ as its input. To address continuous domain shifts during test time, an AdaBN layer is incorporated into the adaptive module. This layer updates its mean and variance based on batch statistics during test time. The predicted $\delta x$ is then added to the original prediction made by $f_{\theta}$ to compensate for inaccurate predictions caused by operating condition domain shift.

### v2

# TARD: Test-time Domain Adaptation for Robust Fault Detection under Evolving Operating Conditions

@@ Abstract
Fault detection is essential in complex industrial systems to prevent failures and optimize performance by distinguishing abnormal from normal operating conditions. With the growing availability of condition monitoring data, data-driven approaches have increasingly applied in detecting system faults. However, these methods typically require large, diverse, and representative training datasets that capture the full range of operating scenarios, an assumption rarely met in practice, particularly in the early stages of deployment. Industrial systems often operate under highly variable and evolving conditions, making it difficult to collect comprehensive training data. This variability results in a distribution shift between training and testing data, as future operating conditions may diverge from those previously observed ones. Such domain shifts hinder the generalization of traditional models, limiting their ability to transfer knowledge across time and system instances, ultimately leading to performance degradation in practical deployments. To address these challenges, we propose a novel method for continuous test-time domain adaptation, designed to support robust early-stage fault detection in the presence of domain shifts and limited representativeness of training data. Our proposed framework –Test-time domain Adaptation for Robust fault Detection (TARD) – explicitly separates input features into system parameters and sensor measurements. It employs a dedicated domain adaptation module to adapt to each input type using different strategies, enabling more targeted and effective adaptation to evolving operating conditions. We validate our approach on two real-world case studies from multi-phase flow facilities, delivering substantial improvements in both fault detection accuracy and model robustness over existing domain adaptation methods under real-world variability.
@@
This work proposes TARD, a novel continuous test-time adaptation framework for fault detection under evolving operating conditions and limited data.
@@ 1 Introduction
To address this data scarcity, one strategy is to leverage information from ”experienced” fleet units – systems with extensive operational histories and rich datasets. A viable solution is to directly apply models trained on these well-instrumented units to less-monitored or newly deployed ones. This cross-unit deployment aims to enhance fault detection performance across the fleet, particularly for units with limited training data. However, such model transfer is hindered by a fundamental limitation: data-driven models commonly assume that training and testing data are independently and identically distributed (i.i.d) [6]. In reality, industrial systems operate under diverse and dynamically changing conditions, violating this assumption. As a result, significant distribution shifts often exist between different units in the fleet. A model trained on one unit may therefore fail to generalize to another, frequently resulting in high false alarm rates and reduced detection accuracy. These limitations ultimately undermine the ability to fully capitalize on shared fleet knowledge.
@@
Domain adaptation (DA) has been extensively explored to address distribution shifts in data-driven models [7, 8]. These methods aim to align a labeled source domain with a related, unlabeled target domain. However, fault detection in industrial systems presents unique challenges. First, DA approaches typically assume the availability of labeled data in the source domain, which is especially scarce or unavailable under faulty conditions in either the source domain or the target domain. As a result, many commonly applied unsupervised DA methods – which rely on source domain labels – are not applicable to fault detection tasks in industrial settings. Second, most DA methods treat domain shift as a single, static distributional difference between discrete domains. However, in complex industrial systems, the operating conditions evolve continuously, leading to domain shifts even within the same unit over time. Therefore, domain adaptation should go beyond discrete domain transfers and be capable of handling continuously evolving shifts. Relying on a fixed, discrete target domain is insufficient for ensuring robust fault detection in such dynamic environments.
@@
To address these challenges, we propose Test-time domain Adaptation for Robust fault Detection (TARD), a novel framework for unsupervised, continual test-time domain adaptation for fault detection that operates entirely without labeled data in the source domain. TARD enables robust fault detection across different units in industrial systems by continuously adapting to dynamically changing operating conditions during test time. The key novelty of our approach lies in:
  • 1.
    A continuous test-time adaptation strategy specifically designed for the unique challenges of unsupervised fault detection, where adapting to unlabeled test batches that potentially contain faulty samples.
  • 2.
    The explicit separation of input variables into control parameters and sensor measurements, with dedicated adaptation strategies for each. This allows the model to distinguish between changes in operating conditions (reflected in control variables) and system faults (reflected in sensor measurements), significantly reducing false alarms.
  • 3.
    A practical framework that requires only a limited number of normal samples from the target system, making it suitable for early-stage deployment and fleet-wide knowledge transfer.
@@ 2.1 Fault Detection
Given this limitation, the common approach is to first shift the objective to detecting any form of abnormal behavior in the system. This process begins with unsupervised anomaly detection, which involves learning a model of normal system operation and then identifying any data patterns that deviates from the baseline [10]. These detected anomalies indicate potential problems, which can be analyzed further to determine if they represent a fault that could lead to a failure of the system [11, 9, 12].
These unsupervised learning methods can be broadly categorized into three main categories: probabilistic models, one-class classification methods, and reconstruction-based approaches [13]. Probabilistic models aim to approximate the probability distribution of normal data, using the estimated density or an associated score function to assign anomaly scores, where lower values typically indicate more anomalous behavior. Various probabilistic models have been applied for this purpose, ranging from classic density estimation models (e.g. gaussian mixture models) [14] to energy-based models (EBMs) [15], predicting anomalies through estimation of the normal data probability distribution. One-class classification models take a different approach by learning a discriminative decision boundary that encloses the region of normal data, without explicitly estimating the full data distribution [13, 16, 17, 18]. These models aim to learn a compact representation of normal behavior and flag data samples that significantly deviate from the normal data distribution as an anomaly[19, 20]. Reconstruction-based methods, such as autoencoders (AEs), are trained to accurately reconstruct normal data samples. Anomalies are then identified based on reconstruction errors –when a test input deviates significantly from the learned normal patterns, the model struggles to reconstruct it, resulting in a high error that signals a potential fault [21, 22]. These models are expected to fit the data distribution under healthy conditions and then raise an alarm for predictions with large deviations when the test data distribution is significantly different from the learned distribution. Despite their usefulness, these unsupervised approaches require large, diverse, and representative datasets that comprehensively cover normal operating conditions, which is rarely fulfilled during the early operational phase of newly deployed systems, making robust early fault detection particularly challenging.
@@ 2.3 Domain Adaptation Applied to Fault Detection
However, most existing DA techniques rely on strong assumptions. They typically presume that abundant source and target samples are available to capture the underlying data distributions. Mean discrepancy-based methods [34] require concurrent access to both abundant source and target domain data for distribution alignment, along with labeled samples for training regularization. Adversarial methods such as Domain Adversarial Neural Networks (DANN) [37] adapt to target distributions through domain discrimination between source and target domains, yet simultaneously demand labeled source data as training supervision. This assumption rarely holds in industrial settings, particularly for newly installed systems where only limited data can be collected, impeding timely condition monitoring [35]. Access to source data may also be restricted due to privacy or proprietary concerns, making unsupervised domain adaptation unfeasible [1].
Methods such as the one proposed in [35] adopt a one-class classification approach to address the challenge of limited data samples. However, they are generally designed for adaptation to one or more discrete, static target domains. In reality, industrial systems often operate under continuously evolving conditions due to environmental fluctuations, component aging, and ongoing system upgrades, leading to non-stationary distributions that static DA models cannot adequately address [38]. This limitation is particularly evident in applications such as energy systems and manufacturing processes, where operating conditions change continuously rather than shifting between discrete states. As a result, domain adaptation strategies tailored for discrete or static domains are ill-suited to dynamic industrial environments, where continuous, on-the-fly adaptation to domain shifts is required.
Recent advances in test-time adaptation (TTA) have emerged as a promising direction to address the limitations of conventional DA approaches. TTA methods adapt pre-trained models during inference by leveraging only the distribution of the current test batch distribution [38]. Notable TTA methods include Tent [39], which updates batch normalization statistics during inference, and SHOT [40], which combines information maximization with self-supervised pseudo-labeling. By dynamically updating model statistics or minimizing prediction entropy at inference, TTA enables real-time adaptation to distribution shifts, making it particularly attractive for PHM applications in environments with continuously evolving operating conditions [41].
@@
Existing methods tend to address isolated aspects of these challenges, but rarely in combination. Fleet-based approaches enable newly deployed units to leverage fleet-wide knowledge, but often fail to adapt to the significant variability of operating conditions within the fleet itself. While DA methods have gained traction, most cannot handle continuous domain shifts and depend on access to sufficiently representative source and target data. Moreover, TTA in the context of unsupervised detection risks model overfitting due to the absence of supervised faulty data. To the best of our knowledge, there is currently no approach that enables unsupervised fault detection under dynamically changing, data-scarce environments. In summary, prior research has made progress in addressing data scarcity and adaptation challenges for fault detection, but several important gaps remain. Our method is designed to fill these gaps through the following distinct contributions:
  • 1.
    Minimal Data Requirements: Most conventional UDA approaches rely on substantial target domain data and often require labeled source data for regularization. However, in many industrial settings, only limited normal operational data is available from the target domain and faulty data is absent. Our method adapts effectively under these realistic constraints, requiring neither large datasets nor labeled faults.
  • 2.
    Dynamic Operational Environments: Existing methods typically assume static domain characteristics, limiting their robustness in real-world applications where operating conditions evolve over time. Our approach explicitly addresses this gap by continuously adapting to non-stationary distribution shifts in industrial systems.
  • 3.
    Reliable Adaptation in Unsupervised Detection: Existing test-time adaptation methods can incorrectly incorporate fault patterns as part of normal system behavior in unsupervised contexts. Our approach overcomes this limitation through the explicit separation and independent adaptation of control variables and sensor measurements, thereby enabling robust fault detection under domain shifts.
@@ 3.1 Problem Definition
The primary motivation for this research is to enable the transfer of operating experience across systems within a fleet – specifically, from systems with abundant monitoring data to other systems operating while enabling adaptation ability to evolving operating conditions. A key challenge addressed in this work is the complete absence of labeled fault data in both the source and target domains. We assume that only healthy operating data is available and that neither the source nor target systems have experienced any faults during the training period. Additionally, we focus on target systems that are newly taken into operation, where only a limited number of observations can be collected for training. Moreover, the data distribution for these systems can evolve continuously due to changes in operating conditions and environmental factors. The objective is to adapt a prediction model trained on the original system so that it can provide accurate predictions for new systems and fleets, even when only a small amount of training data is available. Given:
@@ 3.2 System Variables
In industrial systems, input data typically comprise a mixture of variables reflecting different aspects of system behavior. To effectively handle domain shifts while preserving anomaly detection performance, it is important to distinguish between variables that represent changes in operating conditions and those that reflect the operational state of the system. Therefore, we define two groups within the input parameters: $X=[x,w]$. $w$ denotes control variables, indicating variables that control system conditions. These variables are set by the operators or by the control system to optimize the performance under specified conditions. $x$ represents sensor measurements, which are sensor signals monitoring system components and reflecting real-time system states. Here, we assume that changes in the distribution of control variables do not necessarily indicate an abnormal status but rather distinct operating conditions to which we should adapt.
@@ 3.3 Reconstruction-based Anomaly Detection Framework
We deploy the paradigm of reconstruction-based pipeline for unsupervised anomaly detection as depicted in Figure 1. This method assumes that a model trained exclusively on normal (healthy) data can learn to reconstruct typical system behavior and deviations from this behavior with large reconstruction errors can be interpreted as potential anomalies. We deploy an autoencoder (AE), denoted as $f_{\theta}$, as our reconstruction model. It is trained exclusively on normal source data samples $X_{s}$ to accurately model the normal data distribution with accurately predicted signal values $\hat{X_{s}}$. The training objective is to minimize the mean-squared error (MSE) between the original data samples $X_{s}$, and their reconstructed counterparts $\hat{X_{s}}$:
@@ 4 Methodology
Figure 2: Test-time domain Adaptation for Robust fault Detection (TARD). a) We first pretrain the reconstruction-based anomaly detection model on the source dataset; b) For domain adaptative prediction, we add a domain adaptive module and train it on the target training data.
@@ 4.2 Test-time Domain Adaptation Anomaly Detection
The core of the TARD framework is its ability to adapt to changing operating conditions during inference. As discussed in Section 3.2, control variables $w$ represent operational inputs that are manually or automatically adjusted during system operation. These variables indicate intentional changes in the system’s configuration and are not themselves indicators of abnormal behavior. Sensor measurements $x$, in contrast, reflect the real-time state of the system and are the primary source for fault detection. Therefore, as shown in Figure 2(a), for the reconstruction-based anomaly detection model, control variables are included alongside sensor measurements as part of the input to the autoencoder while the reconstruction target consists only of the sensor measurements $x$. By excluding control variables from the reconstruction target and subsequent anomaly scoring, the model reduces the likelihood of false alarms triggered by expected changes in operating conditions.
@@
Based on the modified reconstruction framework introduced, we integrate an additional adaptive module, $h_{\phi}$, for test-time domain adaptation to bridge the domain gap between the source and target domains. The decision to incorporate a separate adaptive module, $h_{\phi}$, rather than embedding adaptive layers directly into the reconstruction model, stems from limitations observed in unsupervised anomaly detection. TTA methods, such as AdaBN, adapt to each batch during test time, inevitably fitting the distribution of abnormal data points. This impairs the model’s ability to distinguish between normal and abnormal samples. Instead, our adaptive module takes the predicted value and original controlled system variables as inputs. By excluding monitoring data signals from adaptation, this approach prevents overfitting to the potentially faulty data distribution, thereby preserving the model’s capability to accurately distinguish anomalies.
@@
During the adaptation phase, the pre-trained autoencoder $f_{\theta}$ is frozen, and the adaptive module $h_{\phi}$ is trained on a few target data samples to predict $\Delta x$, aimed at compensating for the large prediction errors due to the domain gap between source and target data. This adaptive module exclusively processes the control variables $w$ as its input. To address continuous domain shifts during test time, an AdaBN layer is incorporated into the adaptive module. This layer updates its mean and variance based on batch statistics during test time. The predicted $\Delta x$ is then added to the original prediction made by $f_{\theta}$ to compensate for inaccurate predictions caused by operating condition domain shift. The training procedure for the adaptive module is detailed in Algorithm 1.
Algorithm 1 Fine-tune Adaptive Module on Target Domain Data
0: Pretrained autoencoder $f_{\theta}$ on source domain data, target domain data $X_{t}$, number of epochs $E_{adapt}$, learning rate $\eta_{adapt}$
1: Initialize adaptive module $h_{\phi}$ with random weights
2: Freeze weights of $f_{\theta}$
3: Initialize optimizer (e.g., Adam) with learning rate $\eta_{adapt}$
4:for epoch =1=1 to $E_{adapt}$ do
5:  for each batch $B$ in $X_{t}$ do
6:   Forward pass through autoencoder: $\hat{B}=f_{\theta}(B)$
7:   Forward pass through adaptive module: $\Delta x=h_{\phi}(B_{control})$
8:   Adjust reconstruction: $\hat{B}_{adjusted}=\hat{B}+\Delta x$
9:   Compute reconstruction loss: $\mathcal{L}_{MSE}=\frac{1}{|B|}\sum_{i=1}^{|B|}(B_{i}-\hat{B}_{adjusted,i})^{2}$
10:   Backward pass: Update $h_{\phi}$ using $\nabla\mathcal{L}_{MSE}$
11:  end for
12:end for
13:
14:return Fine-tuned adaptive module $h_{\phi}$